# Tradeoffs of Diagonal Fisher Information Matrix Estimators

### v2

@@ Appendix B Off-Diagonal Variance
We consider an off-diagonal version of the bound given by Theorem 4.1. Notice that in terms of the dependence on neural network weights, the only change is splitting the “responsibility” of the $i$th and $j$th parameter norms.
Theorem B.1.
@@
The proof follows similarly to Appendices I and J, where the primary difference is just swapping the regular eigenvalue-like quantities with the $\gamma$ variational forms. ∎
@@
Lemma B.2.
$\gamma_{\max}(A)=s_{\max}(A)$, where $s_{\max}(A)$ is the maximum singualar value of $A$.
@@
Appendix C Bounding the Trace Variance by Full Spectrum
Theorem C.1.
@@ Appendix C Bounding the Trace Variance by Full Spectrum
The proof follows from a generalized Ruhe’s trace inequality (Marshall et al., 2011):
Theorem C.2.
@@
Thus, noting that the eigenvalue of the “squared” matrix is the matrix’s singular value $\lambda_{t}(J^{\top}J)=s^{2}_{t}(J)$, with Theorem C.2, we have that:
@@
Thus, again simplifying the eigenvalue of the “squared” matrix, with Theorem C.2, we have that:
@@
thus $\mathrm{dHes}^{ia}(\bm{h}\,|\,\bm{x})=\partial^{2}_{ii}(\bm{h}_{a}\,|\,\bm{x})$.
${\mathcal{V}}_{2}(\bm{\theta}\,|\,\bm{x})$ $=\frac{1}{N}\sum_{i=1}^{\dim(\bm{\theta})}\partial^{2}_{ii}{\bm{h}}^{a}(\bm{x})\partial^{2}_{ii}{\bm{h}}^{b}(\bm{x})\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})$
$=\frac{1}{N}\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})\sum_{i=1}^{\dim(\bm{\theta})}\partial^{2}_{ii}{\bm{h}}^{a}(\bm{x})\partial^{2}_{ii}{\bm{h}}^{b}(\bm{x})$
@@
Thus, again simplifying the eigenvalue of the “squared” matrix, with Theorem C.2, we have that:
@@
Appendix D Second Central Moment of Categorical Distribution
@@
Appendix E Empirical Results Continued
In the following section we present additional details and results for the experimental verification we conduct in Section 5.3.
E.I Additional Details
We note that to calculate the diagonal Hessians required for the bounds and empirical FIM calculations, we utilize the BackPACK (Dangel et al., 2019) for PyTorch. Additionally, to calculate the sufficient statistics moment’s spectrum, we explicitly solve the minimum and maximum eigenvalues via their optimization problems. For 2D tensors / matrices, we utilize numpy.linalg.eig. For 4D tensors, we utilize PyTorch Minimize (Feinman, 2021), a wrapper for SciPy’s optimize function.
E.II Additional Plots
We present Figs. I, II, III and IV which are the exact same experiment run in Section 5.3, but with different initial NN weights and random inputs.
Figures VII, VII, VII, VIII and IX show the experimental results on a 5-layer MLP and log-sigmoid activation function. In most of the cases, the FIM and its associated variances quickly go to zero in the first few epochs.
Figure I: The Fisher information, its variances and bounds of the variances w.r.t. a MLP trained with different initialization and a different input $\bm{x}$ (a)
Figure II: The Fisher information, its variances and bounds of the variances w.r.t. a MLP trained with different initialization and a different input $\bm{x}$ (b)
Figure III: The Fisher information, its variances and bounds of the variances w.r.t. a MLP trained with different initialization and a different input $\bm{x}$ (c)
Figure IV: The Fisher information, its variances and bounds of the variances w.r.t. a MLP trained with different initialization and a different input $\bm{x}$ (d)
Figure V: The Fisher information, its variances and bounds of the variances w.r.t. a 5-layer MLP with log-sigmoid activation.
Figure VI: The Fisher information, its variances and bounds of the variances w.r.t. a 5-layer MLP with log-sigmoid activation.
Figure VII: The Fisher information, its variances and bounds of the variances w.r.t. a 5-layer MLP with log-sigmoid activation.
Figure VIII: The Fisher information, its variances and bounds of the variances w.r.t. a 5-layer MLP with log-sigmoid activation.
@@ E.II Additional Plots
Appendix F “Empirical Fisher” Continued
Noting Lemma 6.1’s characterization of the covariance, we are able to characterize the variance of the diagonal elements of $\hat{\mathrm{I}}(\bm{\theta}\,|\,\bm{x})$, denoted as ${\mathrm{V}}(\theta_{i}\,|\,\bm{x})\vcentcolon=\mathrm{Var}(\hat{\mathrm{I}}(\theta_{i}\,|\,\bm{x}))$.
Corollary F.1.
@@
As a result of the similarity of the functional forms of the empirical Fisher $\hat{\mathrm{I}}(\bm{\theta})$ and the FIM estimator ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$, it is not surprising that Corollary F.1 is similar to the variance of ^$\hat{\mathcal{I}}_{1}(\theta_{i}\,|\,\bm{x})$. Indeed, applying Lemma 6.1 will give the exact same functional form with the 2nd central moments of $\bm{t}(\bm{y})$ w.r.t. $p(\bm{y}\,|\,\bm{x})$ exchanged with 2nd non-central moments of $(\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x}))$ w.r.t. $q(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})$. ${\mathrm{V}}(\theta_{i}\,|\,\bm{x})$ is therefore determined by the 2nd and the 4th moment of $(\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x}))$ up to the parameter transformation $\bm{\theta}\to\bm{h}$. Subsequently, the bounds presented for ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ (Eqs. 11 and 4.6) can be similarly adapted for ${\mathrm{V}}(\theta_{i}\,|\,\bm{x})$.
@@
Corollary F.2.
${\mathrm{V}}(\theta_{i})$ $=\frac{1}{N}\mathrm{Var}\left(\mathrm{I}(\theta_{i}\,|\,\bm{x})\right)+\mathop{\mathbb{E}}_{q(\bm{x})}\left[{\mathrm{V}}(\theta_{i}\,|\,\bm{x})\right],$ (32)
@@
If $q(\bm{x},\bm{y})=\frac{1}{N}\sum_{k=1}^{N}\delta(\bm{x}-\bm{x}_{k})\cdot\delta(\bm{y}-\hat{\bm{y}}_{k})$ for a set of observations $\{(\bm{x}_{k},\hat{\bm{y}}_{k})\}_{k=1}^{N}$, then one can directly evaluate the DFIM without sampling and achieve zero variance, i.e., $\hat{\mathrm{I}}(\bm{\theta})=\mathrm{I}(\bm{\theta})$. In this scenario, there is a clear trade-off between the estimators of the FIM in Eqs. 3 and 4 and the DFIM. The estimators of the FIM are unbiased, but have a variance; while the DFIM has zero variance, but is a biased approximation of the FIM.
Appendix G Derivation of Eqs. 5 and 6
@@
where we note that $F^{\prime}_{a}(\bm{h}(\bm{x}_{k}))=\bm{\eta}_{a}(\bm{x}_{k})$ which follows from the connection to expected parameters and partition functions of exponential families, see e.g. Soen & Sun (2021).
@@
The proof follows immediately from the Courant-Fischer min-max theorem Tao (2012). That is,
@@
We note that this can be similarly proven via trace bounds, e.g., Wang et al. (1986). ∎
@@
The proof follows from Lemma A.1, Eq. 23, and directly applying Lemma H.1. ∎
Appendix I Proof of Eq. 11
@@
The proof follows similarly to Lemma H.1. We simple use the following calculation:
@@
From Lemma A.1, we have that
@@
which follows directly from Lemma I.1
We now bound the second term in a similar way, taking $v^{a}\vcentcolon=\partial_{i}{\bm{h}}^{a}(\bm{x})$ and noting that
@@
which follows from Lemma H.1.
@@
Appendix J Proof of Eq. 12
From Lemma A.1 we have that,
${\mathcal{V}}_{2}^{i}=\frac{1}{N}\partial^{2}_{ii}{\bm{h}}^{a}(\bm{x})\partial^{2}_{ii}{\bm{h}}^{b}(\bm{x})\mathcal{I}_{ab}(\bm{h}_{L}).$
Thus, short-handing $h^{a}\vcentcolon=\partial^{2}_{ii}{\bm{h}}^{a}(\bm{x})$, we get
$\|h\|^{2}_{2}\cdot\lambda_{\min}\left(\mathcal{I}(\bm{h})\right)\leq h^{a}h^{b}\cdot\mathcal{I}_{ab}(\bm{h})\leq\|h\|^{2}_{2}\cdot\lambda_{\max}\left(\mathcal{I}(\bm{h})\right),$
which follows from Lemma H.1. This immediately gives the bound as required.
Appendix K Proof of Corollary 4.2
@@
Appendix L Proof of Proposition 4.3
@@
$\lambda_{\max}\left(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})\right)$ $=\sup_{\bm{u}:\|\bm{u}\|_{2}=1}{\bm{u}}^{a}{\bm{u}}^{b}{\bm{u}}^{c}{\bm{u}}^{d}\mathcal{K}^{p}_{abcd}(\bm{t}\,|\,\bm{x})$
@@
Appendix M Proof of Corollary 4.6
@@
Taking the appropriate $\alpha$ and $\beta$ from Eqs. 11 and 12 proves the case for Eqs. 17 and 19.
@@
$\sqrt{\sum_{i=1}^{\dim(\bm{\theta})}\left(\|\partial^{2}_{ii}{\bm{h}}(\bm{x})\|^{2}_{2}\right)^{2}}$ $=\sqrt{\sum_{i=1}^{\dim(\bm{\theta})}\left(\sum_{t=1}^{T}\left[\partial^{2}_{ii}{h_{t}}(\bm{x})\right]^{2}\right)^{2}}$
$\leq\sqrt{\left(\sum_{i=1}^{\dim(\bm{\theta})}\sum_{t=1}^{T}\left[\partial^{2}_{ii}{h_{t}}(\bm{x})\right]^{2}\right)^{2}}$
$=\sum_{i=1}^{\dim(\bm{\theta})}\sum_{t=1}^{T}\left[\partial^{2}_{ii}{h_{t}}(\bm{x})\right]^{2}$
@@
Appendix N Proof of Theorem 4.7
The proof follows from the law of total variances directly. However, we also show how to directly prove the theorem for each estimator individually.
@@
Let $\bm{\delta}_{a}(\bm{x},\bm{y})\vcentcolon=(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))$.
@@
Appendix O Proof of Lemma 4.8
@@
$=\mathop{\mathbb{E}}_{q(\bm{x})}\left[\mathcal{I}(\theta_{i}\,|\,\bm{x})^{2}\right]$
$=\mathop{\mathbb{E}}_{q(\bm{x})}\left[\|\partial_{i}\bm{h}(\bm{x})\|_{2}^{4}\cdot\lambda^{2}_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))\right].$
Appendix P Proof of Proposition 5.1
We first derive the statistics $\mathcal{I}(\bm{h}\,|\,\bm{x})$ and $\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})$ presented in Section 5.1. It follows that from the regression setting, we have that,
@@
Furthermore, by (Soen & Sun, 2021, Lemma 5), we have that,
@@
Appendix Q Proof of Theorem 5.2
@@
Now by using Corollaries 4.2 and Q.1, the remainder of the proof, all we require is the bounding of $\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))$.
@@
Appendix R Proof of Lemma 6.1
@@
The proof follows from the standard definition of covariance. Denoting $\hat{\bm{\eta}}(\bm{x})\vcentcolon=\mathop{\mathbb{E}}_{q(\bm{y}\,|\,\bm{x})}\left[\bm{t}(\bm{y})\right]$, we have:
@@
Appendix S Proof of Corollary F.1
@@
Each of the terms can be calculated: Let $\bm{\delta}_{a}(\bm{x},\bm{y})\vcentcolon=(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))$.
@@
Appendix T Proof of Corollary F.2

### v3

@@ Appendix C Off-Diagonal Variance
We consider an off-diagonal version of the bound given by Theorem 4.1. Notice that in terms of the dependence on neural network weights, the only change is splitting the “responsibility” of the $i$’th and $j$’th parameter norms.
Theorem C.1.
@@
The proof follows similarly to Appendices J and K, where the primary difference is just swapping the regular eigenvalue-like quantities with the $\gamma$ variational forms. ∎
@@
Lemma C.2.
$\gamma_{\max}(A)=s_{\max}(A)$, where $s_{\max}(A)$ is the maximum singular value of $A$.
@@
Appendix D Bounding the Trace Variance by Full Spectrum
Theorem D.1.
@@ Appendix D Bounding the Trace Variance by Full Spectrum
The proof follows from a generalized Ruhe’s trace inequality [Marshall et al., 2011]:
Theorem D.2.
@@
Thus, noting that the eigenvalue of the “squared” matrix is the matrix’s singular value $\lambda_{t}(J^{\top}J)=s^{2}_{t}(J)$, with Theorem D.2, we have that:
@@
Thus, again simplifying the eigenvalue of the “squared” matrix, with Theorem D.2, we have that:
@@
thus $\mathrm{dHes}^{ia}(\bm{h}\,|\,\bm{x})=\partial^{2}_{i}(\bm{h}_{a}\,|\,\bm{x})$.
${\mathcal{V}}_{2}(\bm{\theta}\,|\,\bm{x})$ $=\frac{1}{N}\sum_{i=1}^{\dim(\bm{\theta})}\partial^{2}_{i}{\bm{h}}^{a}(\bm{x})\partial^{2}_{i}{\bm{h}}^{b}(\bm{x})\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})$
$=\frac{1}{N}\mathcal{I}_{ab}(\bm{h}\,|\,\bm{x})\sum_{i=1}^{\dim(\bm{\theta})}\partial^{2}_{i}{\bm{h}}^{a}(\bm{x})\partial^{2}_{i}{\bm{h}}^{b}(\bm{x})$
@@
Thus, again simplifying the eigenvalue of the “squared” matrix, with Theorem D.2, we have that:
@@
Appendix E Second Central Moment of Categorical Distribution
@@
Appendix F Empirical Results Continued
In the following section we present additional details and results for the experimental verification we conduct in Section 5.
F.I Additional Details
We note that to calculate the diagonal Hessians required for the bounds and empirical FIM calculations, we utilize the BackPACK [Dangel et al., 2020] for PyTorch. Additionally, to calculate the sufficient statistics moment’s spectrum, we explicitly solve the minimum and maximum eigenvalues via their optimization problems. For 2D tensors / matrices, we utilize numpy.linalg.eig. For 4D tensors, we utilize PyTorch Minimize [Feinman, 2021], a wrapper for SciPy’s optimize function.
F.II Additional Plots
We present Figs. V, VI, VII and VIII which are the exact same experiment run in Section 5, but with different initial NN weights and random inputs.
Figures XI, XI, XI, XII and XIII show the experimental results on a 5-layer MLP and log-sigmoid activation function. In most of the cases, the FIM and its associated variances quickly go to zero in the first few epochs.
Figure V: The Fisher information, its variances and bounds of the variances w.r.t. a MLP trained with different initialization and a different input $\bm{x}$ (a)
Figure VI: The Fisher information, its variances and bounds of the variances w.r.t. a MLP trained with different initialization and a different input $\bm{x}$ (b)
Figure VII: The Fisher information, its variances and bounds of the variances w.r.t. a MLP trained with different initialization and a different input $\bm{x}$ (c)
Figure VIII: The Fisher information, its variances and bounds of the variances w.r.t. a MLP trained with different initialization and a different input $\bm{x}$ (d)
@@ F.II Additional Plots
Figure X: The Fisher information, its variances and bounds of the variances w.r.t. a 5-layer MLP with log-sigmoid activation.
Figure XI: The Fisher information, its variances and bounds of the variances w.r.t. a 5-layer MLP with log-sigmoid activation.
Figure XII: The Fisher information, its variances and bounds of the variances w.r.t. a 5-layer MLP with log-sigmoid activation.
Figure XIII: The Fisher information, its variances and bounds of the variances w.r.t. a 5-layer MLP with log-sigmoid activation.
Appendix G “Empirical Fisher” Continued
Noting Lemma 6.1’s characterization of the covariance, we are able to characterize the variance of the diagonal elements of $\hat{\mathrm{I}}(\bm{\theta}\,|\,\bm{x})$, denoted as ${\mathrm{V}}(\theta_{i}\,|\,\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathrm{Var}(\hat{\mathrm{I}}(\theta_{i}\,|\,\bm{x}))$.
Corollary G.1.
@@
As a result of the similarity of the functional forms of the empirical Fisher $\hat{\mathrm{I}}(\bm{\theta})$ and the FIM estimator ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$, it is not surprising that Corollary G.1 is similar to the variance of ^$\hat{\mathcal{I}}_{1}(\theta_{i}\,|\,\bm{x})$. Indeed, applying Lemma 6.1 will give the exact same functional form with the 2nd central moments of $\bm{t}(\bm{y})$ w.r.t. $p(\bm{y}\,|\,\bm{x})$ exchanged with 2nd non-central moments of $(\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x}))$ w.r.t. $q(\bm{\hat{\mathrm{y}}}\,|\,\bm{x})$. ${\mathrm{V}}(\theta_{i}\,|\,\bm{x})$ is therefore determined by the 2nd and the 4th moment of $(\bm{t}(\bm{\hat{\mathrm{y}}})-\bm{\eta}(\bm{x}))$ up to the parameter transformation $\bm{\theta}\to\bm{h}$. Subsequently, the bounds presented for ${\mathcal{V}}_{1}(\theta_{i}\,|\,\bm{x})$ (Eqs. 8 and 4.6) can be similarly adapted for ${\mathrm{V}}(\theta_{i}\,|\,\bm{x})$.
@@
Corollary G.2.
Given $N_{x}$ samples of $\bm{x}\sim q(\bm{x})$ and $N$ samples of $\bm{y}_{\,|\,\bm{x}}\sim q(\bm{y}\,|\,\bm{x})$ for each $\bm{x}$ sampled,
${\mathrm{V}}(\theta_{i})=\frac{1}{N_{x}}\cdot\mathrm{Var}\left(\mathrm{I}(\theta_{i}\,|\,\bm{x})\right)+\frac{1}{N_{x}}\cdot\mathop{\mathbb{E}}_{q(\bm{x})}\left[{\mathrm{V}}(\theta_{i}\,|\,\bm{x})\right].$ (25)
@@
If $q(\bm{x},\bm{y})=\frac{1}{N}\sum_{k=1}^{N}\delta(\bm{x}-\bm{x}_{k})\cdot\delta(\bm{y}-\hat{\bm{y}}_{k})$ for a set of observations $\{(\bm{x}_{k},\hat{\bm{y}}_{k})\}_{k=1}^{N}$, then one can directly evaluate the DFIM without sampling and achieve zero variance, i.e., $\hat{\mathrm{I}}(\bm{\theta})=\mathrm{I}(\bm{\theta})$. In this scenario, there is a clear trade-off between the estimators of the FIM in Eq. 3 and the DFIM. The estimators of the FIM are unbiased, but have a variance; while the DFIM has zero variance, but is a biased approximation of the FIM.
Appendix H Derivation of Eq. 3 Using Log-Partition Function Derivatives
In what follows, we derive the alternative equations for ^$\hat{\mathcal{I}}_{1}(\theta_{i})$ and ^$\hat{\mathcal{I}}_{2}(\theta_{i})$ presented in Section 3. That is, we seek to derive the following equations:
^$\hat{\mathcal{I}}_{1}(\theta_{i})$ $=\frac{1}{N}\sum_{k=1}^{N}\left(\frac{\partial F(\bm{h}(\bm{x}_{k}))}{\partial\theta_{i}}-\frac{\partial\bm{h}^{a}(\bm{x}_{k})}{\partial\theta_{i}}\cdot\bm{t}_{a}(\bm{y}_{k})\right)^{2};$ (26)
@@
where we note that $F^{\prime}_{a}(\bm{h}(\bm{x}_{k}))=\bm{\eta}_{a}(\bm{x}_{k})$ which follows from the connection to expected parameters and partition functions of exponential families, see e.g. Soen and Sun [2021].
@@
The proof follows immediately from the Courant-Fischer min-max theorem Tao [2012]. That is,
@@
We note that this can be similarly proven via trace bounds, e.g., Wang et al. [1986]. ∎
@@
The proof follows from Lemma 3.1, Eq. 4, and directly applying Lemma I.1. ∎
Appendix J Proof of Eq. 8
@@
The proof follows similarly to Lemma I.1. We simple use the following calculation:
@@
From Lemma 3.1, we have that
@@
which follows directly from Lemma J.1
We now bound the second term in a similar way, taking $v^{a}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\partial_{i}{\bm{h}}^{a}(\bm{x})$ and noting that
@@
which follows from Lemma I.1.
@@
Appendix K Proof of Eq. 9
From Lemma 3.1 we have that,
${\mathcal{V}}_{2}^{i}=\frac{1}{N}\partial^{2}_{i}{\bm{h}}^{a}(\bm{x})\partial^{2}_{i}{\bm{h}}^{b}(\bm{x})\mathcal{I}_{ab}(\bm{h}_{L}).$
Thus we get
$\|\partial^{2}_{i}{\bm{h}}(\bm{x})\|^{2}_{2}\cdot\lambda_{\min}\left(\mathcal{I}(\bm{h})\right)\leq\partial^{2}_{i}{\bm{h}}^{b}(\bm{x})\cdot\partial^{2}_{i}{\bm{h}}^{b}(\bm{x})\cdot\mathcal{I}_{ab}(\bm{h})\leq\|\partial^{2}_{i}{\bm{h}}(\bm{x})\|^{2}_{2}\cdot\lambda_{\max}\left(\mathcal{I}(\bm{h})\right),$
which follows from Lemma I.1. This immediately gives the bound as required.
Appendix L Proof of Corollary 4.2
@@
Appendix M Proof of Proposition 4.3
@@
$\tilde{\lambda}_{\max}\left(\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})\right)$ $=\sup_{\bm{u}:\|\bm{u}\|_{2}=1}{\bm{u}}^{a}{\bm{u}}^{b}{\bm{u}}^{c}{\bm{u}}^{d}\mathcal{K}^{p}_{abcd}(\bm{t}\,|\,\bm{x})$
@@
Appendix N Proof of Corollary 4.6
@@
Taking the appropriate $\alpha$ and $\beta$ from Eqs. 8 and 9 proves the case for Eqs. 13 and 15.
@@
$\sqrt{\sum_{i=1}^{\dim(\bm{\theta})}\left(\|\partial^{2}_{i}{\bm{h}}(\bm{x})\|^{2}_{2}\right)^{2}}$ $=\sqrt{\sum_{i=1}^{\dim(\bm{\theta})}\left(\sum_{t=1}^{T}\left[\partial^{2}_{i}{h_{t}}(\bm{x})\right]^{2}\right)^{2}}$
$\leq\sqrt{\left(\sum_{i=1}^{\dim(\bm{\theta})}\sum_{t=1}^{T}\left[\partial^{2}_{i}{h_{t}}(\bm{x})\right]^{2}\right)^{2}}$
$=\sum_{i=1}^{\dim(\bm{\theta})}\sum_{t=1}^{T}\left[\partial^{2}_{i}{h_{t}}(\bm{x})\right]^{2}$
@@
Appendix O Proof of Theorem 4.7
To prove the Theorem, we will utilize the law of total variances. We note, that by the premise of the Theorem, we are sampling $N_{x}$ many samples from $q({\bm{x}})$ and $N$ many samples from $q({\bm{y}}\,|\,{\bm{y}})$ for each ${\bm{y}}$ initially sampled. To make this clear, the samples and sampling will be notated by:
${\bm{x}}_{k}$ $\sim q({\bm{x}})$
${\bm{y}}_{l\,|\,{\bm{x}}_{k}}$ $\sim p({\bm{y}}\,|\,{\bm{x}}_{k})$
Note that using these samples, our empirical estimators for the FIM (for either estimator) will be of the form:
^$\hat{\mathcal{I}}_{1}(\theta_{i})=\frac{1}{N_{x}}\sum_{{\bm{x}}_{k}}\left(\frac{1}{N}\sum_{{\bm{y}}_{l\,|\,{\bm{x}}_{k}}}f({\bm{x}}_{k},{\bm{y}}_{l\,|\,{\bm{x}}_{k}})\right),$
for an appropriately chosen $f$.
This also gives:
^$\hat{\mathcal{I}}_{j}(\theta_{i}\,|\,{\bm{x}})=\frac{1}{N}\sum_{{\bm{y}}_{l\,|\,{\bm{x}}}}f({\bm{x}},{\bm{y}}_{l\,|\,{\bm{x}}}).$
Now, we simplify the variance as follows:
$\mathrm{Var}\left[\frac{1}{N_{x}}\sum_{{\bm{x}}_{k}}\left(\frac{1}{N}\sum_{{\bm{y}}_{l\,|\,{\bm{x}}_{k}}}f({\bm{x}}_{k},{\bm{y}}_{l\,|\,{\bm{x}}_{k}})\right)\right]$
$=\frac{1}{N_{x}^{2}}\sum_{{\bm{x}}_{k}}\left(\mathrm{Var}\left[\frac{1}{N}\sum_{{\bm{y}}_{l\,|\,{\bm{x}}_{k}}}f({\bm{x}}_{k},{\bm{y}}_{l\,|\,{\bm{x}}_{k}})\right]\right)$
$=\frac{1}{N_{x}^{2}}\sum_{{\bm{x}}_{k}}\mathrm{Var}\left[\hat{\mathcal{I}}_{j}(\theta_{i}\,|\,{\bm{x}}_{k})\right]$
$=\frac{1}{N_{x}^{2}}\sum_{{\bm{x}}_{k}}\left(\mathrm{Var}_{{\bm{x}}_{k}}\left[\mathop{\mathbb{E}}_{{\bm{y}}_{1\,|\,{\bm{x}}_{k}},\ldots,{\bm{y}}_{N\,|\,{\bm{x}}_{k}}}\left[\hat{\mathcal{I}}_{j}(\theta_{i}\,|\,{\bm{x}}_{k})\right]\right]+\mathop{\mathbb{E}}_{{\bm{x}}_{k}}\left[\mathrm{Var}_{{\bm{y}}_{1\,|\,{\bm{x}}_{k}},\ldots,{\bm{y}}_{N\,|\,{\bm{x}}_{k}}}\left[\hat{\mathcal{I}}_{j}(\theta_{i}\,|\,{\bm{x}}_{k})\right]\right]\right)$
$=\frac{1}{N_{x}^{2}}\sum_{{\bm{x}}_{k}}\left(\mathrm{Var}_{{\bm{x}}_{k}}\left[\mathcal{I}(\theta_{i}\,|\,{\bm{x}}_{k})\right]+\mathop{\mathbb{E}}_{{\bm{x}}_{k}}\left[{\mathcal{V}}_{j}(\theta_{i}\,|\,{\bm{x}}_{k})\right]\right)$
$=\frac{1}{N_{x}}\left(\mathrm{Var}_{{\bm{x}}}\left[\mathcal{I}(\theta_{i}\,|\,{\bm{x}})\right]+\mathop{\mathbb{E}}_{{\bm{x}}}\left[{\mathcal{V}}_{j}(\theta_{i}\,|\,{\bm{x}})\right]\right).$
As required.
@@
Let $\bm{\delta}_{a}(\bm{x},\bm{y})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))$.
@@
Appendix P Proof of Lemma 4.8
@@
$\leq\mathop{\mathbb{E}}_{q(\bm{x})}\left[\mathcal{I}(\theta_{i}\,|\,\bm{x})^{2}\right]$
$\leq\mathop{\mathbb{E}}_{q(\bm{x})}\left[\|\partial_{i}\bm{h}(\bm{x})\|_{2}^{4}\cdot\lambda^{2}_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))\right].$
Appendix Q Proof of Proposition 5.1
We first derive the statistics $\mathcal{I}(\bm{h}\,|\,\bm{x})$ and $\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})$ presented in “Regression: Isotropic Gaussian Distribution” Section 5. It follows that from the regression setting, we have that,
@@
Furthermore, by [Soen and Sun, 2021, Lemma 5], we have that,
@@
Appendix R Proof of Theorem 5.2
@@
Now by using Corollaries 4.2 and R.1, the remainder of the proof, all we require is the bounding of $\lambda_{\max}(\mathcal{I}(\bm{h}\,|\,\bm{x}))$.
@@
Appendix S Proof of Lemma 6.1
@@
The proof follows from the standard definition of covariance. Denoting $\hat{\bm{\eta}}(\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathop{\mathbb{E}}_{q(\bm{y}\,|\,\bm{x})}\left[\bm{t}(\bm{y})\right]$, we have:
@@
Appendix T Proof of Corollary G.1
@@
Each of the terms can be calculated: Let $\bm{\delta}_{a}(\bm{x},\bm{y})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))$.
@@
Appendix U Proof of Corollary G.2